OpenStudy (anonymous):

For the functions f(x) = 2x + 3 and g(x) = 6x + 2, which composition produces the greatest output? Neither composition produces an output. Both compositions produce the same output. f(g(x)) produces the greatest output.<----MY Answer g(f(x)) produces the greatest output.

OpenStudy (anonymous):

help anyone

OpenStudy (perl):

I would start by evaluating f(g(x)) and g(f(x)) f(g(x)) = f(6x+2) = 2(6x+2) + 3 = 12x + 4 + 3 = 12x + 7 can you evaluate g(f(x)) ?

OpenStudy (anonymous):

no sir. i need help with everything

OpenStudy (anonymous):

im completely lost

OpenStudy (anonymous):

@perl

OpenStudy (lulubj):

I would think its C or D.

OpenStudy (anonymous):

me to i just cant decide

OpenStudy (anonymous):

@neonmeanings

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I think its D